# A Hybrid Supervised and Self-Supervised Graph Neural Network for Edge-Centric Applications

### v1

@@ IV-A Protein-Protein Interaction (PPI) prediction
The experiment was conducted using a 5-fold cross-validation, repeated 10 times, to mitigate the effect of random initialization. This procedure averaged the results over multiple runs, reducing variability in the final performance metrics. The same folds were used by our model and the baseline for a fair comparison. Preliminary experiments indicate that a tokenizer composed of two linear layers using ReLU and a final Tanh activation function produced the best results.
Figure 4: Comparison of the performance between our model and the SVGAE model across different organisms. Each boxplot represents the distribution of the F1 scores obtained from the 10 times 5-fold cross-validation. The boxplots show the mean F1 score for each k-fold, and the small box inside each plot indicates the overall mean F1 score across all k-folds. The p-value from the Wilcoxon signed-rank test is also indicated, showing the significance level of the difference between models.
To evaluate the performance of the proposed model, we compared it with the Signed Variational Graph Auto-Encoder (SVGAE) [22], a state-of-the-art approach specifically designed for protein-protein interaction (PPI) prediction. SVGAE has demonstrated significant accuracy improvements over traditional sequence-based methods and achieved leading performance on various datasets described in the corresponding section. SVGAE was selected as a benchmark due to its effective utilization of graph structures for capturing complex relationships in biological networks, aligning well with the edge-centric design of their model.
Figure 4 shows the F1 score performance of both models on each dataset. Each boxplot represents the distribution of F1 scores from the 10 repeated runs of 5-fold cross-validation, with boxplot pairs comparing our model to SVGAE on each dataset. As indicated by the Wilcoxon signed-rank test, our model significantly outperformed SVGAE on four datasets—C. elegans, Drosophila, E. coli, and Human—with notably low p-values (p $<$ 0.001). For the HPRD dataset, however, there was no significant difference in performance between the two models, suggesting comparable predictive accuracy in this specific case.
@@ IV-B Gene Ontology (GO) term prediction
Figure 5: The critical difference (CD) diagram presents the statistical significance of the results. Models connected by a black line have no statistically significant difference. The CD value is 1.36.
The Friedman test results indicate that there are significant differences in the performance of the models being compared (p=$1e\!-\!6$). The critical difference diagram in Figure 5 shows that both the proposed model and exp2GO are the best methods for gene function prediction, with no statistically significant difference between them (p = $0.237$). However, there is a significant difference between the proposed model and deepGOplus (p = $31e\!-\!$ 3) as well as NMFGO (p = $31e\!-\!$ 3).
@@ IV-C Compounds similarity prediction
Figure 7: Comparison of embedding spatial distribution using MSE (right) and combined cost function (left). The combined cost function improves the linear relationship between compound similarity and cosine distance.
Figure 7 shows two subplots comparing predicted similarity (y-axis) to the cosine distance of embeddings (x-axis) under different cost functions. The right subplot displays results using our combined cost function, while the left shows the Mean Squared Error (MSE) approach. The combined cost function not only enhances the spatial organization of embeddings but also establishes a clearer linear relationship between compound similarity and cosine distance, facilitating accurate predictions. This arrangement enables both compounds with known and unknown structures to be effectively ordered within the embedding space, optimizing the predictor’s performance.
@@ V Conclusion
In conclusion, the proposed model offers a flexible and powerful solution for a range of graph-based applications. Future research may focus on extending its capabilities to other fields of application, for example in recommendation systems, and social network analytics, among others that may benefit from the capabilities of predicting properties of pairs of nodes. In all our experiments we have used a one-layer MPNN, but we need to explore the possibility of using additional layers of MPNN in the Embedder section. Furthermore, future investigations will consider methods to generate predictions with associated uncertainty intervals, providing a measure of confidence in the model’s predictions.

### v2

@@ IV-A Protein-Protein Interaction (PPI) prediction
The experiment was conducted using a 5-fold cross-validation, repeated 10 times, to mitigate the effect of random initialization. This procedure averaged the results over multiple runs, reducing variability in the final performance metrics. The same folds were used by our model and the baseline for a fair comparison. Preliminary experiments indicate that an input tokenizer composed of two linear layers using ReLU and a final Tanh activation function produced the best results.
Figure 5: Critical-difference (CD) diagram of average ranks for the PPI experiments. With a CD value of 1.83 ($\alpha=0.05$), no horizontal black lines appear, meaning every pair of models differs by more than the CD and thus all performance differences are statistically significant.
To evaluate the performance of the proposed model, we compared it against several state-of-the-art approaches. First, we included the Signed Variational Graph Auto-Encoder (SVGAE) [18], a method specifically designed for PPI tasks that uses graph structures to capture complex relationships in biological networks, achieving superior performance compared to traditional sequence-based models. Additionally, we considered the Siamese Residual RCNN model proposed by Chen et al. [31], a deep learning approach that applies residual connections and convolutional architectures to predict PPIs from sequence information. Finally, we evaluated EnAmDNN (Ensemble Deep Neural Networks with Attention Mechanism) [32], the leading sequence-based predictor reported in the recent benchmark by Dunham et al. [33], which demonstrated remarkable accuracy by integrating ensemble learning and attention mechanisms.
Figure 4 summarises the F1-score distributions for the PPI task. For each of the five organisms, four side-by-side boxplots compare our model with SVGAE, Siamese Residual RCNN, and EnAmDNN. Every boxplot reflects the scores obtained in 10 repetitions of 5-fold cross-validation; the line inside the box denotes the mean F1 of each k-fold.
The Friedman test reports a significant overall difference among the four PPI models ($p=7.8e\!-\!30$). The critical-difference diagram in Figure 5 displays no connecting lines, indicating that every pair of methods differs by more than the critical difference. Our model achieves the highest average rank, with the Siamese Residual RCNN in second place; the gap between them is statistically significant ($p=1.5e\!-\!2$).
@@ IV-B Gene Ontology (GO) term prediction
Figure 7: The critical difference (CD) diagram presents the statistical significance of the results. Models connected by a black line have no statistically significant difference. The CD value is 1.36 ($\alpha=0.05$).
The Friedman test results indicate that there are significant differences in the performance of the models being compared ($p=1e\!-\!6$). The critical difference diagram in Figure 7 shows that both the proposed model and exp2GO are the best methods for gene function prediction, with no statistically significant difference between them ($p=0.237$). However, there is a significant difference between the proposed model and deepGOplus ($p=1e\!-\!3$) as well as NMFGO ($p=1e\!-\!3$).
@@ IV-C Compounds similarity prediction
Ablation study on compound similarity
To better understand the individual contributions of the loss components and the attention mechanism, we performed an ablation study using the compound similarity dataset. Two independent experiments were carried out in addition to a basic baseline MPNN. The attention type experiment kept the composite loss $L_{\mathrm{s}}+L_{\mathrm{c}}+L_{\mathrm{cp}}$ fixed and evaluated four attention configurations: none, node only, edge only, and the combined node&edge scheme. Conversely, the loss type experiment held the node&edge attention constant and assessed four loss formulations: $L_{\mathrm{s}}$, $L_{\mathrm{s}}+L_{\mathrm{cp}}$, $L_{\mathrm{s}}+L_{\mathrm{c}}$, and the full $L_{\mathrm{s}}+L_{\mathrm{c}}+L_{\mathrm{cp}}$.
Figure 10: Mean Absolute Error (MAE) for the three models—baseline MPNN, attention experiment and loss experiment—presented in that order. Lower values indicate better performance.
Figure 11: Predicted compound similarity (y-axis) versus cosine distance between embeddings (x-axis) under four ablation settings: baseline model without attention and self-supervised losses (top-left), attention only (top-right), hybrid loss only (bottom-left), and the full model with both attention and hybrid loss (bottom-right).
Figure 10 displays boxplots summarising the Mean Absolute Error obtained in each of the five folds of cross-validation for the baseline model and for every configuration evaluated in the attention-type and loss-type experiments. Each box therefore represents the distribution of errors across the 5-fold splits, providing a concise view of the variability and central tendency associated with each setting.
Figure 11 plots, for each compound pair, the cosine distance between embeddings on the horizontal axis and the similarity predicted by the model on the vertical axis. The four panels correspond to the baseline configuration, the model with attention only, the model with the hybrid loss only, and the full model. In the baseline panel, the points form a wide, diffuse cloud, indicating that embedding distance is not a good proxy for similarity prediction. Enabling either component separately narrows the cloud and makes the trend more clearly monotonic, showing that each element independently helps the embeddings capture similarity. When both the attention mechanism and the hybrid losses are active, the points collapse into an almost straight, tight band; the markedly lower dispersion indicates that the embeddings are better organised in the latent space, thereby enabling the model to predict similarity much more accurately. This improved organisation not only accounts for the lower MAE observed but also ensures that compounds lacking structural information are embedded coherently alongside known compounds, providing a reliable latent representation for downstream analyses.
@@ V Conclusion
In conclusion, the proposed model offers a flexible and powerful solution for a wide range of graph-based applications, particularly for tasks that require evaluating the properties of pairs of objects. Future research may focus on extending its capabilities to other fields of application, for example in recommendation systems, and social network analytics, among others that may benefit from the capabilities of predicting properties of pairs of nodes. In all our experiments we have used a one-layer MPNN, but we need to explore the possibility of using additional layers of MPNN in the Embedder section. Furthermore, future investigations will consider methods to generate predictions with associated uncertainty intervals, providing a measure of confidence in the model predictions.